HanAkoSolo (jamierox4ev3r):

y=x^2-2x-15 This problem wants me to find the vertex, the x-and y-intercepts, and the symmetric point. Then I have to graph. help plz? >.<

HanAkoSolo (jamierox4ev3r):

I have an idea on how to go about solving this…pretty sure it involves the quadratic formula but idk so yeah. ._.

OpenStudy (nincompoop):

convert it into vertex form y = a(x-h)^2 + k where (h, k) is your vertex

OpenStudy (nincompoop):

this is usually achieved by completing the square

HanAkoSolo (jamierox4ev3r):

oh that thing ._. wouldn't it be like: y+15=x^2-2x y+15+___=x^2-2x+____ y+15+___=(x+___)^2

HanAkoSolo (jamierox4ev3r):

that thing?

HanAkoSolo (jamierox4ev3r):

EW!

OpenStudy (nincompoop):

I don't know why you moved 15 all the way to the other side of the equation, but it should work as long as you put it back where it belongs

OpenStudy (zale101):

take half of the x coefficient and then square it

OpenStudy (nincompoop):

y=x^2-2x-15 y = (x^2 - 2x + ______) - ____ - 15 y = (x^2 - 2x + 1) - 1 - 15 y= x^2 -2x + 1 is a complete square so you can now simplify it by factorization then process the -1 -15 this is now your k

OpenStudy (nincompoop):

half of the x of the middle term though, because x^2 is also an x LOL @Zale101

OpenStudy (nincompoop):

so if we were to continue @Jamierox4ev3r we end up with y = (x-1)^2 +16

HanAkoSolo (jamierox4ev3r):

wait wut how?

HanAkoSolo (jamierox4ev3r):

hold on… ugh history homework overload in process

OpenStudy (nincompoop):

do one subject at a time

HanAkoSolo (jamierox4ev3r):

no i mean i just finished my history…and it kinda gave me a headache ._.

OpenStudy (nincompoop):

y+15=x^2-2x y+15+___=x^2-2x+____ y+15+___=(x+___)^2 we can work with your way of solving it after finding the last term to make it a perfect square, we need to add the same thing on the left hand side of the equation, correct?

HanAkoSolo (jamierox4ev3r):

oh wait…ha i just figured it out c;

HanAkoSolo (jamierox4ev3r):

ha right all the blanks would be filled with 1..

HanAkoSolo (jamierox4ev3r):

and you would subtract 16 from both sides, thus getting y=(x-1)^2+16

OpenStudy (nincompoop):

because the half of the middle term would be -2/2 = -1 then square (-1)^2 = 1 so y + 15 + 1 = x^2 -2x + 1

HanAkoSolo (jamierox4ev3r):

yeah i see now…figuring that out was like a catharsis xDD

HanAkoSolo (jamierox4ev3r):

so that means that the vertex point is (-1,16) right? and the y-intercept would be (0,15)

HanAkoSolo (jamierox4ev3r):

still need to find the x-intercepts tho, and that I do with the quadratic formula

OpenStudy (nincompoop):

y + 16 = (x-1)^2 then add (-16) on both sides of the equation y + 16 + (-16) = (x-1)^2 + (-16) y = (x-1)^2 - 16 to look for the x-intercept, your y value will be zero so 0 = (x-1)^2 - 16

HanAkoSolo (jamierox4ev3r):

okay i follow so far

OpenStudy (nincompoop):

solve for x

HanAkoSolo (jamierox4ev3r):

wait quick question tho… would my vertex point and y-intercept be correct?

OpenStudy (nincompoop):

y intercept means your x = 0

HanAkoSolo (jamierox4ev3r):

right…and thats what i have up there ^ i think…*checks* yeah it is

HanAkoSolo (jamierox4ev3r):

oh wait, and i realized i can plug in the original equation into my calculator and check if everything is correct :P alright then

HanAkoSolo (jamierox4ev3r):

oh wait…its (1,-16)

OpenStudy (nincompoop):

ftw

OpenStudy (nincompoop):

did you figure out your x intercept ?

HanAkoSolo (jamierox4ev3r):

uh yeah..working on it

OpenStudy (nincompoop):

0 = (x-1)^2 - 16 16 = (x-1)^2 square root each side 4 = x - 1 and -4 = x -1

HanAkoSolo (jamierox4ev3r):

wait… oh mi gawsh thats so … ._. wow i was over thinking this 4=x-1 -4=x-1 +1 +1 +1 +1 ----- ------- x=5 x=-3

HanAkoSolo (jamierox4ev3r):

so the x-intercepts are (5,0) and (-3,0)

OpenStudy (nincompoop):

you were probably using the quadratic formula?

HanAkoSolo (jamierox4ev3r):

mhm

OpenStudy (nincompoop):

try solving it using that method

HanAkoSolo (jamierox4ev3r):

wow holy *____fill in the blank____* wish my math teach taught us that ._. but then again, Mr. Tom Sawyer never explains anything

OpenStudy (nincompoop):

Write the quadratic polynomial on the left hand side in standard form. Expand out terms of the left hand side: x^2-2 x+1 = 16 Move everything to the left hand side. Subtract 16 from both sides: x^2-2 x-15 = 0 Using the quadratic formula, solve for x. x = (2±sqrt((-2)^2-4 (-15)))/2 = (2±sqrt(4+60))/2 = (2±sqrt(64))/2: x = (2+sqrt(64))/2 or x = (2-sqrt(64))/2 Simplify radicals. sqrt(64) = sqrt(2^6) = 8 = 8: x = (2+8)/2 or x = (2-8)/2 Evaluate (2+8)/2. (2+8)/2 = 10/2 = 5: x = 5 or x = (2-8)/2 Evaluate (2-8)/2. (2-8)/2 = -6/2 = -3: Answer: | | x = 5 or x = -3

HanAkoSolo (jamierox4ev3r):

that is so much easier!!!

OpenStudy (nincompoop):

which is?

HanAkoSolo (jamierox4ev3r):

this way, off *mind explosion*

HanAkoSolo (jamierox4ev3r):

way easier ofc :P

OpenStudy (nincompoop):

both are easy

HanAkoSolo (jamierox4ev3r):

stupid auto correct gawsh! but thanks so much wow I'm so grateful for the help, this one makes more sense to me for some reason

OpenStudy (nincompoop):

it is just a matter of knowing which method to use given a particular situation

HanAkoSolo (jamierox4ev3r):

and to graph it…all i have to do is use the info (x-ints, y int., vertex)

OpenStudy (nincompoop):

exactly

OpenStudy (nincompoop):

you'll do this even in college, so you better learn it now

HanAkoSolo (jamierox4ev3r):

aha…EHEHEHEHEHEHEHEHE I GET IT NOW ._. okay I'm calm now I swear, but seriously thank you so much :P

OpenStudy (nincompoop):

how about this nifty solution y=x^2-2x-15 vertex (-b/2a, y)

HanAkoSolo (jamierox4ev3r):

wait okay okay so uh what am i looking at here?

HanAkoSolo (jamierox4ev3r):

just, you know, some quick clarification >_>

OpenStudy (nincompoop):

a = 1, b = -2, c = -15 vertex (-b/2a, y) -b/2a -(-2)/2(1) 2/2 = 1 y = x^2 - 2x - 15, where x = 1 y = 1^2 - 2(1) -15 y = 1 - 2 - 15 y = -16

HanAkoSolo (jamierox4ev3r):

oh. well what do you know…i see ^_^

OpenStudy (nincompoop):

this method does not rely on completing the square, but it involves you to know that the x in vertex is solved by using the formula -b/2a no completing the square, no moving things around and no factoring involved

HanAkoSolo (jamierox4ev3r):

i see…..have you ever considered being a math teacher? LOL anyhow i see how this works now, i prefer to avoid completing the square and stuff like that

OpenStudy (nincompoop):

oh yeah definitely works very well and saves you a lot of time and when it means fewer steps, fewer mistakes could be made. it does not let you off the hook solving the x and y intercepts

HanAkoSolo (jamierox4ev3r):

oh yeah, but that's fine :) less possibilities for mistakes are always nifty